mhchen:

Linear Algebra: Find Coefficients a,b,c,d such that the equation: ax^2 + ay^2 + bx + cy + d = 0 runs through the points: (-2,7) (-4,5) (4,-3)

mhchen:

What I did so far: Create a matrix representing [a b c d] for each point: \[\left[\begin{matrix}53 & -2 & 7 & 1 \\ 41 & -4 & 5 & 1 \\ 25 & 4 & -3 & 1 \end{matrix}\right]\] Then I did some matrix operations until I arrive at this point: \[\left[\begin{matrix}1 & 0 & * & * \\ 0 & 1 & * & * \\ 0 & 0 & 1 & *\end{matrix}\right]\] Where * represents a real number. But at this point, I think I noticed a problem. The 4th column (which represents the coefficient of 'd', will always be a real-number instead of 1 or 0. Do you guys see the same problem?

mhchen:

I think I did my matrix wrong from the very beginning

xXMarcelieXx:

have you checked it on symbolab for matrix ?

Narad:

I got the augmented matrix as \[\left[\begin{matrix}4 & 49&5&1\\ 16 & 25&1&1\\16&9&16&1\end{matrix}\right]\]

mhchen:

@Narad I think there's an issue with your matrix. The first row: [4 49 5 1] seems to correspond to the point (-2,7) right? ax^2 = -2^2 = 4 ay&2 = 7^2 = 49 but for the rest: bx = -2 cy = 7 d = 1 So I'm not seeing how you made your first row. I know the matrix is consisted of coefficients for a,b,c,d: (x^2 + y^2)a + (x)b + (y)c + (1)d = 0 \[\left[\begin{matrix}x^2+y^2 & x & y & 1 & | & 0\end{matrix}\right]\]

mhchen:

That's how I got the matrix: \[\left[\begin{matrix}(-2)^2+(7)^2 & -2 & 7 & 1 & | & 0\\ (-4)^2+(5)^2 & -4 & 5 & 1 & | & 0 \\ (4)^2+(-3)^2 & 4 & -3 & 1 & | & 0\end{matrix}\right]\] Row 1 is for (-2,7), x = -2, y = 7 Row 2 is for (-4,5), x = -4, y = 5 Row 3 is for (4,-3), x = 4, y = -3 which becomes \[\left[\begin{matrix}53 & -2 & 7 & 1 & | & 0 \\ 41 & -4 & 5 & 1 & | & 0\\ 25 & 4 & -3 & 1 & | & 0 \end{matrix}\right]\]

mhchen:

@xXMarcelieXx So I took your advice and it turns out I made a mistake lol: https://www.symbolab.com/solver/matrix-calculator/gauss%20jordan%20%5Cbegin%7Bpmatrix%7D53%26-2%267%261%260%5C%5C%2041%26-4%265%261%260%5C%5C%2025%264%26-3%261%260%5Cend%7Bpmatrix%7D Even so I'm not sure how to find the values of a,b,c,d based on this matrix: \[\left[\begin{matrix}1 & 0 & 0 & \frac{1}{29} & | & 0 \\ 0 & 1 & 0 & \frac{-2}{29} & | & 0\\ 0 & 0 & 1 & \frac{-4}{29} & | & 0 \end{matrix}\right]\] Which is solving this system of equation: \[a + \frac{1}{29}d = 0\] \[b + \frac{-2}{29}d = 0\] \[c + \frac{-4}{29}d = 0\] I could set d = r, and make it a parametric equation.

Narad:

Second row of the matrix corresponding to the point (-4,5) a*(-4)^2 + b*(5)^2 +c(-4+5)+d=0 16a+25b+1c+1d=0 The coefficients are 16, 25, 1,1 And you can determine the coefficients of the the third point (4,-3)

mhchen:

But the equation of the circle I'm given is ax^2 + ay^2 + bx + cy + d = 0 not ax^2 + by^2 + c(x+y) + d

mhchen:

I am SO sorry I wrote my question wrong, I corrected it.

mhchen:

WOW PARAMATRIC EQUATIONS WORK LOLWUT: https://www.desmos.com/calculator/rmerfpca9w r can be anything except 0.

xXMarcelieXx:

oh okay i see lol i remember doing this like 4 semesters ago but its pretty easy but i forgot lol

mhchen:

@xXMarcelieXx oh wow so you've taken linear algebra before? Mind giving me some advice on this course, like which topics are the hardest

mhchen:

Also I'm taking Mathematical Analysis which is similar to the questions you ask. (Proving mathematical statements"

xXMarcelieXx:

yeah, i took a class that was combined with diff equations and linear algebra. I dont remmeber much but i think the last section with longer equations were harder. My advice is do flashcards for certain problems and vocab stuff. Math proofs? @mhchen

mhchen:

@xXMarcelieXx yes math proofs. Real numbers, natural numbers, set theory, and I just started cardinality

xXMarcelieXx:

oh, you are way ahead lol im barely on direct proofs, conditonals and biconditonals. Perhaps, do you have any notes ? what math textbook do you use?

mhchen:

I'm in like a level higher than you so the textbook I use is way too hard for you. You can just look at the first page to see if you understand: http://cms.dm.uba.ar/academico/materias/verano2012/taller_de_calculo_avanzado/Libros/Abbott%20-%20Understanding%20Analysis.pdf Yeah I know you're starting with truth tables, and learning about notations.

xXMarcelieXx:

oh thats way too advanced for me lmao

mhchen:

Better for you cause then it's easier for me to help you out.